# Reconstruction of boosted and resolved multi-Higgs-boson events with symmetry-preserving attention networks

### v1

@@ Abstract
The production of multiple Higgs bosons at the CERN LHC provides a direct way to measure the trilinear and quartic Higgs self-interaction strengths as well as potential access to beyond the standard model effects that can enhance production at large transverse momentum $p_{\mathrm{T}}$. The largest event fraction arises from the fully hadronic final state in which every Higgs boson decays to a bottom quark-antiquark pair ($\mathrm{b\overline{b}}$). This introduces a combinatorial challenge known as the jet assignment problem: assigning jets to sets representing Higgs boson candidates. Symmetry-preserving attention networks (SPA-Nets) have been been developed to address this challenge. However, the complexity of jet assignment increases when simultaneously considering both $\mathrm{H}\to\mathrm{b\overline{b}}$ reconstruction possibilities, i.e., two “resolved” small-radius jets each containing a shower initiated by a $\mathrm{b}$ quark or one “boosted” large-radius jet containing a merged shower initiated by a $\mathrm{b\overline{b}}$ pair. The latter improves the reconstruction efficiency at high $p_{\mathrm{T}}$. In this work, we introduce a generalization to the SPA-Net approach to simultaneously consider both boosted and resolved reconstruction possibilities and unambiguously interpret an event as “fully resolved,” “fully boosted,” or in between. We report the performance of baseline methods, the original SPA-Net approach, and our generalized version on nonresonant $\mathrm{HH}$ and $\mathrm{HHH}$ production at the LHC. Considering both boosted and resolved topologies, our SPA-Net approach increases the Higgs boson reconstruction purity by 57–62% and the efficiency by 23–38% compared to the baseline method depending on the final state.
@@ 1 Introduction
Measuring multiple Higgs boson production at the LHC is a powerful probe of the Higgs trilinear ($\lambda_{3}$) and quartic ($\lambda_{4}$) self-couplings as well as new physics effects beyond the standard model (SM), such as resonant production via a heavy particle decaying into two Higgs bosons. While the quartic coupling is challenging to measure at the LHC and requires future colliders to observe the corresponding signature of triple Higgs boson production Brigljevic:2024vuv ; Cepeda:2019klc ; Plehn:2005nk , the trilinear coupling can be directly accessed by studying Higgs boson pair production at the LHC. These processes can also exhibit modified kinematics and enhanced cross sections at large transverse momentum $p_{\mathrm{T}}$ for non-SM coupling values Dolan:2013rja ; Dolan:2015zja ; Bishara:2016kjn ; Carvalho_2016 ; Arganda:2018ftn ; Capozi_2020 . The dominant decay mode of the Higgs boson is into a bottom quark-antiquark ($\mathrm{b\overline{b}}$) pair, which leads to a fully hadronic final state with multiple jets. Representative Feynman diagrams for the production of two or three Higgs bosons in the gluon-gluon fusion process where all Higgs bosons decay to $\mathrm{b\overline{b}}$ pairs are shown in Fig. 1. A key challenge in this final state is to find the optimal association of jets to Higgs boson candidates. This problem is combinatorial in nature and becomes computationally challenging for conventional methods as the number of jets increases.
@@
Moreover, the complexity of this problem is further increased by the presence of different event topologies, depending on the $p_{\mathrm{T}}$ of the Higgs bosons. At low $p_{\mathrm{T}}$, each Higgs boson can be reconstructed as two “resolved” small-radius jets, each containing a $\mathrm{b}$-quark-initiated jet. At high $p_{\mathrm{T}}$, each Higgs boson can be reconstructed from one “boosted” large-radius jet, containing a merged $\mathrm{b\overline{b}}$ pair. For intermediate $p_{\mathrm{T}}$, a mixture of resolved and boosted jets can be present. This is illustrated in Fig. 2 The CMS and ATLAS Collaborations have performed many searches for resonant and non-resonant Higgs boson pair production in boosted CMS:2018vjd ; CMS:2022gjd and resolved CMS:2018qmt ; CMS:2018sxu ; ATLAS:2020jgy ; CMS:2022cpr final states, and their statistical combination ATLAS:2015zug ; ATLAS:2016paq ; ATLAS:2018rnh ; CMS:2022dwd ; ATLAS:2022hwc ; ATLASCombination2024 ; CMS-PAS-HIG-20-011 . The combination of these two channels is challenging because the two data samples need to be statistically independent, but the two selection criteria can often select the same events, especially for Higgs bosons in the intermediate $p_{\mathrm{T}}$ range. Thus, analysts must implement an “overlap removal” procedure to allow these two searches to be combined. Part of the motivation of this work is to use machine learning to help interpret individual events as either boosted or resolved.
@@
Figure 2: Two bottom quarks from a Higgs boson decay are reconstructed as two separated small-$R$ jets in the resolved topology when the Higgs boson $p_{\mathrm{T}}\lesssim 2m_{\mathrm{H}}/R$ (left). Two bottom quarks from a Higgs boson decay overlap and merge into a single large-$R$ jet in the boosted topology when the Higgs boson $p_{\mathrm{T}}\gtrsim 2m_{\mathrm{H}}/R$ (center). An event with two Higgs bosons where one is resolved and one is boosted (right)
@@
Symmetry-preserving attention networks (SPA-Nets) are a novel class of neural networks that have been proposed Shmakov:2021qdz ; Fenton:2020woz ; Fenton:2023ikr to solve the jet assignment problem for a fixed event topology. They have also been studied in the context of resolved nonresonant $\mathrm{HH}$ events to improve the sensitivity to $\lambda_{3}$ at the high-luminosity LHC Chiang:2024pho . In this work, we present a generalization of the SPA-Net approach to handle variable (boosted and resolved) event topologies and simultaneously consider both small- and large-radis jets. We apply this approach to nonresonant $\mathrm{HH}$ and $\mathrm{HHH}$ production at the LHC and compare its performance with baseline methods and the original fully resolved SPA-Net approach. Finally, we also demonstrate how SPA-Net can provide a partitioning of events into disjoint categories based on the number of reconstructed boosted Higgs boson candidates.
@@
The rest of this paper is organized as follows. Section 2 presents the dataset and ground truth labeling. We describe the baseline methods and SPA-Net model configurations in Section 3. In Section 4, we delineate how we use the SPA-Net outputs to assign jets and categorize events. We present the results in Section 5 and summarize in Section 6.
@@
The dataset duarte_2024_14257266 and code javier_duarte_2024_14248961 for analyzing it are publicly available. For implementing the training, we use the SPANet library version 2.2 with additions related to reweighting and an improved validation accuracy metric described in Section 3 in a custom fork on GitHub spanet .
@@ 2 Dataset and Labeling
Nonresonant $\mathrm{HH}$ and $\mathrm{HHH}$ events are generated using MadGraph5_amc@nlo3.4.1 Alwall:2014hca at the LHC with a center of mass energy $\sqrt{s}=14\,\text{TeV}$, with the trilinear and quartic couplings fixed to the SM values as implemented in the model of Ref. Papaefstathiou:2020lyp . Only the decay of the Higgs boson to bottom quarks is considered. The parton shower and hadronization are simulated with pythia8.2 Sjostrand:2014zea and the detector response is simulated with delphes3.4.1 deFavereau:2013fsa , using the CMS parametrization.
Final-state particles are clustered into jets using the anti-$k_{\mathrm{T}}$ algorithm Cacciari:2008gp ; Cacciari:2011ma with radius parameters of 0.5 (AK5 jets) and 0.8 (AK8 jets). We require $p_{\mathrm{T}}>20\,\text{GeV}$ for AK5 jets and $p_{\mathrm{T}}>200\,\text{GeV}$ for AK8 jets, as well as $\lvert\eta\rvert<2.5$ for both. Emulation of a $\mathrm{b}$-tagging algorithm is applied to the AK5 jets, which assigns a boolean value to each jet indicating whether it originates from a $\mathrm{b}$ quark, based on $p_{\mathrm{T}}$-dependent efficiency and misidentification rates. The misidentification rate for light ($\mathrm{c}$) jets is about 1% (20%), while the $\mathrm{b}$-tagging efficiency is about 70% CMS:2012feb . We note that these are conservative relative to current state-of-the-art $\mathrm{b}$-tagging algorithms Mondal:2024nsa . The four-momentum $(p_{\mathrm{T}},\eta,\phi,m)$ and the $\mathrm{b}$-tagging value of each AK5 jet are stored as inputs to the networks. For AK8 jets, we consider the four-momentum $(p_{\mathrm{T}},\eta,\phi,m)$ as well. No $\mathrm{b}$-tagging is applied for AK8 jets, however, we develop our own tagger discussed in Section 3.1.
@@
Figure 3: Reconstruction efficiency for true Higgs bosons to be reconstructed and matched to either two AK5 jets or one AK8 jet as a function of $p_{\mathrm{T}}$ in $\mathrm{HHH}$ events.
@@
We also generated a background dataset from events composed uniquely of jets produced through the strong interaction, referred to as quantum chromodynamics (QCD) multijet events, using the same generator as for the non-resonant $\mathrm{HH}$ and $\mathrm{HHH}$ events. To ensure a sufficient sample size with real high-$p_{\mathrm{T}}$ $\mathrm{b}$ jets, we generated 4 $\mathrm{b}$ partons and required their scalar $p_{\mathrm{T}}$ sum to be $H_{\mathrm{T}}>250\,\text{GeV}$. After applying the same pre-selection criteria as for the $\mathrm{HHH}$ events, the QCD test dataset consists of approximately 25,000 events.
@@
$\mathrm{HHH}$, $m_{\mathrm{H}}=125\,\text{GeV}$ 855,254 90.25% 4.75% 5%
$\mathrm{HHH}$, $m_{\mathrm{H}}\in\{120,122.5,125,127.5,130\}\,\text{GeV}$ 9,580,718 95% 5%
$\mathrm{HH}$, $m_{\mathrm{H}}=125\,\text{GeV}$ 1,209,807 90.25% 4.75% 5%
@@ 3.1 Baseline
where $m(j_{a},j_{b})$ is the invariant mass of a pair of AK5 jets $j_{a}$ and $j_{b}$ and $m_{\mathrm{H}}=125\,\text{GeV}$ is the nominal mass of the Higgs boson. The AK5 jets are sorted in descending order by $p_{\mathrm{T}}$, with the $\mathrm{b}$-tagged sorted jets coming first and the non-$\mathrm{b}$-tagged coming last. For the $\chi^{2}$ method, only the first six jets are considered.
For the boosted topology baseline, the delphes configuration lacks a parameterized boosted Higgs boson AK8 jet tagger in delphes, so we train a boosted decision tree (BDT) tagger as a baseline to compare with SPA-Net. Given an AK8 jet, the BDT input variables are the $p_{\mathrm{T}}$, $\eta$, mass $m$, soft-drop mass $m_{\mathrm{SD}}$ Larkoski:2014wba , fraction of charged energy, number of charged particles, and the $N$-subjettiness jet substructure variables $\tau_{32}$ and $\tau_{21}$ Thaler:2010tr . The BDT output score represents the probability that the AK8 jet is an $\mathrm{H}\to\mathrm{b\overline{b}}$ jet rather than a QCD jet. It was trained on a dataset that combined the AK8 jets in the SM $\mathrm{HHH}$ training set and the AK8 jets in the QCD training set, which consists of 164,601 $\mathrm{H}\to\mathrm{b\overline{b}}$ jets and 553,548 QCD jets. The classes were balanced during training.

### v3

@@ Abstract
The production of multiple Higgs bosons at the CERN LHC provides a direct way to measure the trilinear and quartic Higgs self-interaction strengths as well as potential access to beyond the standard model effects that can enhance production at large transverse momentum $p_{\mathrm{T}}$. The largest event fraction arises from the fully hadronic final state in which every Higgs boson decays to a bottom quark-antiquark pair ($\mathrm{b\overline{b}}$). This introduces a combinatorial challenge known as the jet assignment problem: assigning jets to sets representing Higgs boson candidates. Symmetry-preserving attention networks (SPA-Nets) have been been developed to address this challenge. However, the complexity of jet assignment increases when simultaneously considering both $\mathrm{H}\to\mathrm{b\overline{b}}$ reconstruction possibilities, i.e., two “resolved” small-radius jets each containing a shower initiated by a $\mathrm{b}$ quark or one “boosted” large-radius jet containing a merged shower initiated by a $\mathrm{b\overline{b}}$ pair. The latter improves the reconstruction efficiency at high $p_{\mathrm{T}}$. In this work, we introduce a generalization to the SPA-Net approach to simultaneously consider both boosted and resolved reconstruction possibilities and unambiguously interpret an event as “fully resolved,” “fully boosted,” or in between. We report the performance of baseline methods, the original SPA-Net approach, and our generalized version on nonresonant $\mathrm{HH}$ and $\mathrm{HHH}$ production at the LHC. Considering both boosted and resolved topologies, our SPA-Net approach increases the Higgs boson reconstruction purity by 56–80% and the efficiency by 37–38% compared to the baseline method depending on the final state.
@@ 1 Introduction
Measuring multiple Higgs boson production at the LHC is a powerful probe of the Higgs trilinear ($\lambda_{3}$) and quartic ($\lambda_{4}$) self-couplings as well as new physics effects beyond the standard model (SM), such as resonant production via a heavy particle decaying into two Higgs bosons. While the quartic Higgs self-coupling is notoriously difficult to measure at the LHC, current theoretical studies indicate that even at a future 100  TeV hadron collider the Standard Model $\mathrm{HHH}$ production process would be observable with only a 2–3 $\sigma$ significance Cepeda:2019klc ; Plehn:2005nk ; Papaefstathiou:2019ofh ; Fuks:2017zkg ; Contino:2016spe . Nevertheless, triple Higgs production remains a direct probe of this coupling as well as physics beyond the SM and has recently attracted increasing attention in LHC studies  Brigljevic:2024vuv . By contrast, the trilinear self-coupling can be accessed more directly through measurements of Higgs boson pair production DiMicco:2019ngk . These processes can also exhibit modified kinematics and enhanced cross sections at large transverse momentum ($p_{\mathrm{T}}$) for non-SM coupling values Dolan:2013rja ; Dolan:2015zja ; Bishara:2016kjn ; Carvalho_2016 ; Arganda:2018ftn ; Capozi_2020 , which makes it important to accurately reconstruct and measure them across a wide range of $p_{\mathrm{T}}$. The dominant decay mode of the Higgs boson is into a bottom quark-antiquark ($\mathrm{b\overline{b}}$) pair, which leads to a fully hadronic final state with multiple jets. Representative Feynman diagrams for the production of two or three Higgs bosons in the gluon-gluon fusion process where all Higgs bosons decay to $\mathrm{b\overline{b}}$ pairs are shown in Fig. 1. A key challenge in this final state is to find the optimal association of jets to Higgs boson candidates. This problem is combinatorial in nature and becomes computationally challenging for conventional methods as the number of jets increases.
@@
Moreover, the complexity of this problem is further increased by the presence of different event topologies, depending on the $p_{\mathrm{T}}$ of the Higgs bosons. At low $p_{\mathrm{T}}$, each Higgs boson can be reconstructed as two “resolved” small-radius jets, each containing a $\mathrm{b}$-quark-initiated jet. At high $p_{\mathrm{T}}$, each Higgs boson can be reconstructed from one “boosted” large-radius jet, containing a merged $\mathrm{b\overline{b}}$ pair. For intermediate $p_{\mathrm{T}}$, a mixture of resolved and boosted jets can be present. This is illustrated in Fig. 2. The CMS and ATLAS Collaborations have performed many searches for resonant and nonresonant Higgs boson pair production in boosted CMS:2018vjd ; CMS:2022gjd and resolved CMS:2018qmt ; CMS:2018sxu ; ATLAS:2020jgy ; CMS:2022cpr final states, and their statistical combination ATLAS:2015zug ; ATLAS:2016paq ; ATLAS:2018rnh ; CMS:2022dwd ; ATLAS:2022hwc ; ATLASCombination2024 ; CMS-PAS-HIG-20-011 . In CMS, the boosted category is found to be 40% more sensitive than the resolved, illustrating the importance of having an inclusive reconstruction targeting the various topologies. As of now, there is no experimental result targeting an intermediate event topology reconstruction. In the context of $\mathrm{HH}$ and $\mathrm{HHH}$, maximizing the signal acceptance by targeting the different topologies is important in order to enhance the sensitivity to the SM cross section.
The combination of the boosted and resolved channels is challenging because the two data samples need to be statistically independent, but the two selection criteria can often select the same events, especially for Higgs bosons in the intermediate $p_{\mathrm{T}}$ range. Thus, analysts must implement an “overlap removal” procedure to allow these two searches to be combined. Part of the motivation of this work is to use machine learning to help interpret individual events as boosted, resolved, or in some intermediate configuration.
@@
Figure 2: Two bottom quarks from a Higgs boson decay are reconstructed as two separated small-$R$ jets in the resolved topology when the Higgs boson $p_{\mathrm{T}}\lesssim 2m_{\mathrm{H}}/R$ (left). Here, $m_{\mathrm{H}}$ denotes the Higgs boson mass. Two bottom quarks from a Higgs boson decay overlap and merge into a single large-$R$ jet in the boosted topology when the Higgs boson $p_{\mathrm{T}}\gtrsim 2m_{\mathrm{H}}/R$ (center). An event with two Higgs bosons where one is resolved and one is boosted (right)
@@
Symmetry-preserving attention networks (SPA-Nets) are a novel class of neural networks that have been proposed Shmakov:2021qdz ; Fenton:2020woz ; Fenton:2023ikr to solve the jet assignment problem for a fixed event topology. They have also been studied in the context of resolved nonresonant $\mathrm{HH}$ events to improve the sensitivity to $\lambda_{3}$ at the high-luminosity LHC Chiang:2024pho . In this work, we present a generalization of the SPA-Net approach to handle variable (boosted and resolved) event topologies and simultaneously consider both small- and large-radius jets. We apply this approach to nonresonant $\mathrm{HH}$ and $\mathrm{HHH}$ production at the LHC and compare its performance with baseline methods and the original fully resolved SPA-Net approach. Finally, we also demonstrate how SPA-Net can provide a partitioning of events into disjoint categories based on the number of reconstructed boosted Higgs boson candidates.
@@
The rest of this paper is organized as follows. Section 1.1 discusses related work. Section 2 presents the dataset and ground truth labeling. We describe the baseline methods and SPA-Net model configurations in Section 3. In Section 4, we delineate how we use the SPA-Net outputs to assign jets and categorize events. We present the results in Section 5 and summarize in Section 6.
@@
The dataset duarte_2024_14257266 and code javier_duarte_2024_14248961 for analyzing it are publicly available. For implementing the training, we use the SPANet library version 2.3, which features new additions related to reweighting and an improved validation accuracy metric as described in Section 3 spanet .
1.1 Related Work
Recently, the ATLAS Collaboration has also searched for resonant and nonresonant $\mathrm{HHH}$ production in the $\mathrm{b\overline{b}}\mathrm{b\overline{b}}\mathrm{b\overline{b}}$ final state ATLAS:2024xcs , which finds sensitivity at the level of $750\times SM$ at 95% confidence level. In the context of the search for $\mathrm{HH}\to 4\mathrm{b}$, resolved, intermediate, and boosted topologies were studied in Ref. Amacker:2020bmn where the categorization is defined based on the presence of a large-radius jet. In order to enhance the sensitivity to the $\mathrm{HH}$ process as well as to enable inclusive approaches for searches for the $\mathrm{HHH}$ process at the LHC, we propose a generalization of the jet assignment and event categorization in this paper. Ref. CMS:2022gjd uses a state-of-the-art jet tagger, ParticleNet Qu:2019gqs in order to identify the boosted $\mathrm{H}\to\mathrm{b\overline{b}}$ large-radius jet candidates in the $\mathrm{HH}\to 4\mathrm{b}$ search. Such methods can be used in combination with the method we propose, for example as a training input, making them complimentary. As mentioned above, Ref. Chiang:2024pho applies the standard SPA-Net method with small-radius jets to the resolved nonresonant $\mathrm{HH}\to 4\mathrm{b}$ search and quantifies the potential improvement in the sensitivity to $\lambda_{3}$ at the HL-LHC. While our study does not aim to estimate the full analysis sensitivity as this would require background modeling, systematic uncertainties, and trigger-level considerations that are beyond the scope of our work, we aim to evaluate the improvements in signal reconstruction performance from our proposed enhancement to SPA-Net.
@@ 2 Dataset and Labeling
Nonresonant $\mathrm{HH}$ and $\mathrm{HHH}$ events are generated using MadGraph5_amc@nlo3.4.1 Alwall:2014hca at a leading-order (LO) accuracy at the LHC with a center of mass energy $\sqrt{s}=14\,\text{Te\hskip-0.80002ptV}$, with the trilinear and quartic couplings fixed to the SM values as implemented in the model of Ref. Papaefstathiou:2020lyp . The corresponding $\mathrm{HH}$ and $\mathrm{HHH}$ differential cross sections as a function of the individual $\mathrm{H}$ $p_{\mathrm{T}}$ observables is shown in Fig. 3. Only the decay of the Higgs boson to bottom quarks is considered. The parton shower and hadronization are simulated with pythia8.2 Sjostrand:2014zea and the detector response is simulated with delphes3.4.1 deFavereau:2013fsa , using the CMS parametrization.
Figure 3: $\mathrm{HH}$ (left) and $\mathrm{HHH}$ (right) differential cross sections scaled to next-to-next-to-LO (NNLO) as a function of the individual $\mathrm{H}$ $p_{\mathrm{T}}$ observables. The Higgs bosons are sorted by $p_{\mathrm{T}}$.
Final-state particles are clustered into jets using the anti-$k_{\mathrm{T}}$ algorithm Cacciari:2008gp ; Cacciari:2011ma with radius parameters of 0.5 (AK5 jets) and 0.8 (AK8 jets). We require $p_{\mathrm{T}}>20\,\text{Ge\hskip-0.80002ptV}$ for AK5 jets and $p_{\mathrm{T}}>200\,\text{Ge\hskip-0.80002ptV}$ for AK8 jets, as well as $\lvert\eta\rvert<2.5$ for both. Emulation of a $\mathrm{b}$-tagging algorithm is applied to the AK5 jets, which assigns a boolean value to each jet indicating whether it originates from a $\mathrm{b}$ quark, based on $p_{\mathrm{T}}$-dependent efficiency and misidentification rates. The misidentification rate for light ($\mathrm{c}$) jets is about 1% (20%), while the $\mathrm{b}$-tagging efficiency is about 70% CMS:2012feb . We note that these are conservative relative to current state-of-the-art $\mathrm{b}$-tagging algorithms Mondal:2024nsa . The four-momentum $(p_{\mathrm{T}},\eta,\phi,m)$ and the $\mathrm{b}$-tagging value of each AK5 jet are stored as inputs to the networks. For AK8 jets, we consider the four-momentum $(p_{\mathrm{T}},\eta,\phi,m)$ as well. No $\mathrm{b}$-tagging is applied for AK8 jets, however, we develop our own tagger discussed in Section 3.1.
@@
Figure 4: Reconstruction efficiency for true Higgs bosons to be reconstructed and matched to either two AK5 jets or one AK8 jet as a function of $p_{\mathrm{T}}$ in $\mathrm{HHH}$ events with $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.74ptV}$.
@@
We also generated a background dataset from events composed uniquely of jets produced through the strong interaction, referred to as quantum chromodynamics (QCD) multijet events, using the same generator as for the non-resonant $\mathrm{HH}$ and $\mathrm{HHH}$ events. To ensure a sufficient sample size with real high-$p_{\mathrm{T}}$ $\mathrm{b}$ jets, we generated 4 $\mathrm{b}$ partons and required their scalar $p_{\mathrm{T}}$ sum to be $H_{\mathrm{T}}>250\,\text{Ge\hskip-0.80002ptV}$. After applying the same pre-selection criteria as for the $\mathrm{HHH}$ events, the QCD test dataset consists of approximately 25,000 events.
@@
$\mathrm{HHH}$, $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.80002ptV}$ 855,254 90.25% 4.75% 5%
$\mathrm{HHH}$, $m_{\mathrm{H}}\in\{120,122.5,125,127.5,130\}\,\text{Ge\hskip-0.80002ptV}$ 9,580,718 95% 5%
$\mathrm{HH}$, $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.80002ptV}$ 1,209,807 90.25% 4.75% 5%
@@ 3.1 Baseline
where $m(j_{a},j_{b})$ is the invariant mass of a pair of AK5 jets $j_{a}$ and $j_{b}$ and $m_{\mathrm{H}}=125\,\text{Ge\hskip-0.80002ptV}$ is the nominal mass of the Higgs boson. The AK5 jets are sorted in descending order by $p_{\mathrm{T}}$, with the $\mathrm{b}$-tagged sorted jets coming first and the non-$\mathrm{b}$-tagged coming last. For the $\chi^{2}$ method, only the first six jets are considered.
For the boosted topology baseline, the default delphes configuration lacks a parameterized boosted Higgs boson AK8 jet tagger and does not save the jet constituents for all AK8 jets necessary to train a state-of-the-art jet tagger like ParticleNet Qu:2019gqs or Particle Transformer Qu:2022 . In lieu of that, we train a boosted decision tree (BDT) tagger as a baseline to compare with SPA-Net. Given an AK8 jet, the BDT input variables are the $p_{\mathrm{T}}$, $\eta$, mass $m$, soft-drop mass $m_{\mathrm{SD}}$ Larkoski:2014wba , fraction of charged energy, number of charged particles, and the $N$-subjettiness jet substructure variables $\tau_{32}$ and $\tau_{21}$ Thaler:2010tr . The BDT output score represents the probability that the AK8 jet is an $\mathrm{H}\to\mathrm{b\overline{b}}$ jet rather than a QCD jet. It was trained on a dataset that combined the AK8 jets in the SM $\mathrm{HHH}$ training set and the AK8 jets in the QCD training set, which consists of 164,601 $\mathrm{H}\to\mathrm{b\overline{b}}$ jets and 553,548 QCD jets. The classes were balanced during training. We show the receiver operating characteristic (ROC) curve, qrea under the curve (AUC), and accuracy of the BDT in Fig. 5.
Figure 5: ROC curve for the BDT-based boosted topology baseline, showing the QCD background mistag rate versus the $\mathrm{H}\to\mathrm{b\overline{b}}$ signal efficiency. The working point corresponds to a signal efficiency of 38% and a background misidentification rate of 2%